OpenStudy (swagster23780):

Solve. 4x + 2 ≥ –6 A. x ≤ –1 B. x ≤ –2 C. x ≥ –2 D. x ≥ –1

OpenStudy (onibteram):

The answer is B. 4*-2=-8+2=-6 \[x\ge\] means its either greater than or equal to -6 @SWAGSTER23780